# Multi-Objective min-max Online Convex Optimization

### v1

@@ Appendix F Bounding $R_{3}$ with the adversarial input
Since $\bar{\theta}$ is a particular (constant) element of $\Delta_{2}$, for every $x\in\mathcal{X}$
$\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle\leq\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\langle\theta,\lambda_{t}(x)\rangle,$
and therefore taking the minimum over $x$ yields
$\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle\leq C_{\mathsf{OPT}}.$
Thus,
$R_{3}$ $=\min_{x}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-C_{\mathsf{OPT}}$
$\leq\min_{x}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-\min_{x}\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle$
$\leq\max_{x}\sum_{t=1}^{T}\big\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\big\rangle,$
where the last inequality follows as below.
Let $x^{\star}$ satisfy $u(x^{\star})=\min_{x}u(x)$. Then
$\min_{x}u(x)-\min_{x}v(x)=u(x^{\star})-\min_{x}v(x)\leq u(x^{\star})-v(x^{\star})\leq\max_{x}(u(x)-v(x)),$
which proves the claim.
This proves
$R_{3}\leq\max_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\rangle.$ (29)
Since we are in the bounded setting, for all rounds $t$ and all $x\in\mathcal{X}$ we have $\|\lambda_{t}(x)\|_{\infty}\leq B$. Thus, appyling the Hölder’s inequality (using duality of $||.||_{1}$ and $||.||_{\infty}$ norms) on (29),
$\sum_{t=1}^{T}\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\rangle\leq\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}\,\|\lambda_{t}(x)\|_{\infty}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
Thus,
$R_{3}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
This is the desired inequality. ∎
Next, we state an elementary inequality without proof.
Lemma 12 (Deviation from average via consecutive differences).
Let $\theta_{1},\dots,\theta_{T}\in\mathbb{R}^{d}$ be vectors, and define the average
$\bar{\theta}=\frac{1}{T}\sum_{t=1}^{T}\theta_{t}.$
Then
$\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}\;\leq\;2T\sum_{t=1}^{T-1}\|\theta_{t+1}-\theta_{t}\|_{1}.$
Lemma 13 (One-step TV bound for Hedge).
Let $\{\theta_{t}\}$ be the Hedge distributions updated by
$\theta_{t+1,i}\;=\;\frac{\theta_{t,i}\,e^{\eta_{\theta,t}\lambda_{t,i}}}{\sum_{j}\theta_{t,j}\,e^{\eta_{\theta,t}\lambda_{t,j}}},\qquad i=1,\dots,n,$
with learning rate $\eta_{\theta,t}>0$. Assume the losses satisfy $\lambda_{t,i}\in[0,B]$ for all $i$. Then
$\|\theta_{t+1}-\theta_{t}\|_{1}\leq\eta_{\theta,t}B.$
@@
Let
$Z_{t}:=\sum_{j}\theta_{t,j}e^{\eta_{\theta,t}\lambda_{t,j}}\;=\;\mathbb{E}_{i\sim\theta_{t}}\!\big[e^{\eta_{\theta,t}\lambda_{t,i}}\big].$
By the definition of Kullback–Leibler divergence,
$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})$ $=\sum_{i}\theta_{t,i}\ln\frac{\theta_{t,i}}{\theta_{t+1,i}}=\sum_{i}\theta_{t,i}\ln\frac{\theta_{t,i}}{\theta_{t,i}e^{\eta_{\theta,t}\lambda_{t,i}}/Z_{t}}$
$=\sum_{i}\theta_{t,i}(\ln Z_{t}-\eta_{\theta,t}\lambda_{t,i})=\ln Z_{t}-\eta_{\theta,t}\,\mathbb{E}_{i\sim\theta_{t}}[\lambda_{t,i}].$
Let $\mu:=\mathbb{E}_{i\sim\theta_{t}}[\lambda_{t,i}]$. Then $\ln Z_{t}=\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}\lambda_{t,i}}\big]=\eta_{\theta,t}\mu+\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}\big]$, so (1) is equivalently
$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})=\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}\big].$ (30)
Thus to upper bound the one-step KL it suffices to bound the centered MGF $\ln\mathbb{E}[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}]$.
Proposition 14 (Hoeffding’s MGF bound).
For a random variable $X$ that satisfies $X\in[a,b]$ almost surely and $\mathbb{E}[X]=\mu$, then for any $s\in\mathbb{R}$
$\ln\mathbb{E}[e^{s(X-\mu)}]\leq\frac{s^{2}(b-a)^{2}}{8}.$ (31)
Since $\lambda_{t,i}\in[0,B]$, using (31) in (30) with $X=\lambda_{t,i}$, [a,b]=[B,B][a,b]=[-B,B], $s=\eta_{\theta,t}$, and mean $\mu$, we get
$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})\leq\frac{\eta_{\theta,t}^{2}4B^{2}}{8}.$ (32)
Proposition 15 (Pinsker’s inequality [12]).
For any two probability vectors $p,q$,
$\|p-q\|_{1}\leq\sqrt{2\,D_{\mathrm{KL}}(p\|q)}.$ (33)
Since both $\theta_{t}$ and $\theta_{t+1}$ are distributions, applying (33) on (32), we get
$\|\theta_{t+1}-\theta_{t}\|_{1}\leq\sqrt{2\cdot\frac{\eta_{\theta,t}^{2}B^{2}}{2}}=\eta_{\theta,t}B,$
the claimed bound. ∎

### v3

@@ Appendix I Proof of Theorem 9
The regret decomposition and the bounds on ${\mathbb{E}}\{R_{1}\}$ and ${\mathbb{E}}\{R_{2}\}$ are identical to those in the proof of Theorem 5. It remains to show that ${\mathbb{E}}\{R_{3}\}\leq 0$.
Recall
$R_{3}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-\bar{C}_{\mathsf{OPT}},\qquad\Lambda_{t}(x,\theta)=\theta^{\top}\lambda_{t}(x).$
Let $(x^{\star},\theta^{\star})$ solve
$\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x).$
As in the proof of Lemma 8, we have the deterministic inequality
$R_{3}\leq\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star}).$ (35)
Taking conditional expectation with respect to $\mathcal{F}_{t-1}$ and using (32),
$\mathbb{E}\!\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\mid\mathcal{F}_{t-1}\right\}=(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star}).$
By optimality of $\theta^{\star}$, the right-hand side is non-positive almost surely. Summing over $t$ and taking expectations in (35) yields
${\mathbb{E}}\{R_{3}\}\leq 0.$
Combining Lemma 14 and Theorem 15, the true min-max static regret of Algorithm 1 with respect to the benchmark ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ is $O(\sqrt{T\log T})$ under the Martingale difference condition.
Algorithm 5 Blocked Hedge+OGD (for Markovian Inputs)
1:Input: Block width $H$, step sizes $\eta_{x,m}>0$, $\eta_{\theta,m}>0$, feasible set $\mathcal{X}$
2:Initialize: $w_{1}=(1,1)$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, $x_{1}\in\mathcal{X}$ arbitrary
3:Let $M=\lceil T/H\rceil$ be the number of blocks.
4:for $m=1,\dots,M$ do
5:  1. Freeze: Set current action $x=x_{m}$ and weights $\theta=\theta_{m}$
6:  Set effective block length $H_{m}=\min(H,\,T-(m-1)H)$
7:  2. Initialize block accumulators:
8:  $\bar{g}_{m}=0$, $\bar{\lambda}_{m}=0$
9:  3. Play and Observe (Block Loop):
10:  for $t=(m-1)H+1,\dots,(m-1)H+H_{m}$ do
11:   Play action $x_{t}=x$
12:   Observe losses $(f_{t}(x),g_{t}(x))$ and gradients $(\nabla f_{t}(x),\nabla g_{t}(x))$
13:   Evaluate loss vector $\lambda_{t}=(f_{t}(x),g_{t}(x))$
14:   Form surrogate gradient:
$\nabla\tilde{\Lambda}_{t}(x,\theta):=\theta_{1}\nabla f_{t}(x)+\theta_{2}\nabla g_{t}(x)$
15:   Accumulate:
16:    $\bar{\lambda}_{m}\leftarrow\bar{\lambda}_{m}+\lambda_{t}$
17:    $\bar{g}_{m}\leftarrow\bar{g}_{m}+\nabla\tilde{\Lambda}_{t}(x,\theta)$
18:  end for
19:  4. Block averaging:
20:  $\bar{\lambda}_{m}\leftarrow\bar{\lambda}_{m}/H_{m}$
21:  $\bar{g}_{m}\leftarrow\bar{g}_{m}/H_{m}$
22:  5. Updates (End of Block):
23:  OGD: $y_{m+1}=x_{m}-\eta_{x,m}\bar{g}_{m}$
24:    $x_{m+1}=\mathrm{Proj}_{\mathcal{X}}(y_{m+1})$
25:  Hedge: $w_{m+1,i}=w_{m,i}\exp(\eta_{\theta,m}\bar{\lambda}_{m,i})$, $i=1,2$
26:  Normalize: $\theta_{m+1}=w_{m+1}/\|w_{m+1}\|_{1}$
27:end for
Appendix O Markovian Inputs
In this section, we extend our main result (Theorem 5 of Section 6) to the setting where the loss sequence evolves according to a stationary Markov process. Unlike the Martingale difference case, temporal dependence introduces a persistent bias that requires modifying the algorithm to decouple the learning updates from the instantaneous state of the Markov chain.
Let $\{S_{t}\}_{t\geq 1}$ be a time-homogeneous, ergodic Markov chain on a measurable state space $\mathcal{S}$ with transition kernel $P$ and stationary distribution $\pi$. We assume the chain is uniformly ergodic, i.e., there exist constants $C>0$ and $\tau_{\mathrm{mix}}>0$ such that for any initial distribution $\nu$,
$\|\nu P^{t}-\pi\|_{\mathrm{TV}}\leq Ce^{-t/\tau_{\mathrm{mix}}},\qquad\forall t\geq 1.$
Let $\lambda:\mathcal{X}\times\mathcal{S}\to\mathbb{R}^{2}$ be a measurable mapping and define
$\lambda_{t}(x):=\lambda(x,S_{t})=(f_{t}(x),g_{t}(x)).$
We assume that $f_{t}$ and $g_{t}$ satisfy Assumptions 1 and 2.
Define the stationary mean
$\mu(x):=\mathbb{E}_{\pi}[\lambda(x,S)],\qquad\forall x\in\mathcal{X},$ (36)
which is independent of $t$ by stationarity.
Recall the benchmark
${\mathbb{E}}\{C_{\mathsf{OPT}}\}={\mathbb{E}}\!\left[\min_{x\in\mathcal{X}}\max\!\left\{\sum_{t=1}^{T}f_{t}(x),\;\sum_{t=1}^{T}g_{t}(x)\right\}\right],$ (37)
and the deterministic surrogate benchmark
$\bar{C}_{\mathsf{OPT}}=T\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (38)
and the expected min-max static regret is defined as before
$\bar{R}_{\mathcal{A}}(T)=\sup_{P}\{{\mathbb{E}}\{C_{\mathcal{A}}\}-\bar{C}_{\mathsf{OPT}}\}.$
Lemma 16 (Benchmark Gap under Markovian Inputs).
Under the Markovian model described above and Assumptions 1 and 2,
${\mathbb{E}}\left\{\bigl|C_{\mathsf{OPT}}-\bar{C}_{\mathsf{OPT}}\bigr|\right\}\;\leq\;O(\sqrt{T\log T})+O(\tau_{\mathrm{mix}}).$
@@
The proof follows Appendix D with the sole modification that concentration of $\sum_{t=1}^{T}\lambda_{t}(x)$ around $T\mu(x)$ uses Markov-chain concentration instead of independence. Specifically, using uniform ergodicity and standard additive functional concentration (e.g., [36]), for any fixed $x\in\mathcal{X}$:
$\mathbb{E}\!\left\{\Big\|\sum_{t=1}^{T}\lambda_{t}(x)-T\mu(x)\Big\|_{\infty}\right\}=O(\sqrt{T\log T})+O(\tau_{\mathrm{mix}}).$
All minimax and $\varepsilon$-net arguments of Appendix D remain unchanged, yielding the stated bound. ∎
Standard application of Algorithm 1 fails in the Markovian setting because the algorithm’s updates $\theta_{t}$ become highly correlated with the instantaneous state $S_{t}$, leading to an $O(T)$ regret. To address this, we propose a Blocked Hedge+OGD strategy in Algorithm 5. The time horizon $T$ is divided into $M=\lceil T/H\rceil$ blocks of length $H$. Within block $m$, the algorithm freezes its action $x_{m}$ and weights $\theta_{m}$, accumulates losses, and performs a single update for both the Hedge and the OGD at the end of the block.
Theorem 17.
Under Assumptions 1 and 2, running the Blocked Hedge+OGD algorithm (Algorithm 5) with block size $H=\Theta(T^{1/3}\tau_{\mathrm{mix}}^{2/3})$ yields an expected min-max static regret of
$\bar{R}_{\mathcal{A}}(T)=O(T^{2/3}\tau_{\mathrm{mix}}^{1/3}).$
Proof.
Let $M=T/H$ be the number of blocks. We decompose the regret over the blocks similar to (10). The expected min-max static regret ${\bar{R}}_{\mathcal{A}}(T)$ is decomposed into three terms by adding and subtracting $\sum_{m=1}^{M}\Lambda^{(m)}(x_{m},\theta_{m})$ and $\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\Lambda^{(m)}(x,\theta_{m})$:
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{\mathbb{E}\{R_{1}\}+\mathbb{E}\{R_{2}\}+\mathbb{E}\{R_{3}\}\}$
where:
$R_{1}=\max_{\theta\in\Delta_{K}}\sum_{m=1}^{M}\theta^{\top}\Lambda^{(m)}(x_{m})-\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x_{m}),$
$R_{2}=\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x_{m})-\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x),$
and
$R_{3}=\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x)-{\bar{C}}_{\mathsf{OPT}}.$
In Appendix O.1, we prove the following bounds
$\mathbb{E}\{R_{1}+R_{2}\}\leq O(H\sqrt{M})=O(H\sqrt{T/H})=O(\sqrt{TH}).$ (39)
$\mathbb{E}\{R_{3}\}\leq\sum_{m=1}^{M}O(\tau_{\mathrm{mix}})=\frac{T}{H}O(\tau_{\mathrm{mix}}).$ (40)
Combining these, the expected min-max static regret is bounded by:
$\bar{R}_{\mathcal{A}}(T)\leq O(\sqrt{TH})+O\left(\frac{T\tau_{\mathrm{mix}}}{H}\right).$
Minimizing this bound with respect to $H$ requires balancing $\sqrt{TH}$ with $T\tau_{\mathrm{mix}}/H$, which yields the optimal block length $H=\Theta(T\tau_{\mathrm{mix}}^{2})^{1/3}$. Substituting this back gives the final bound $O(T^{2/3}\tau_{\mathrm{mix}}^{1/3})$. ∎
Combining Lemma 16 and Theorem 17, the true min-max static regret of the Blocked Hedge+OGD Algorithm 5 with respect to the true benchmark ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ is
$O(T^{2/3}\tau_{\mathrm{mix}}).$
O.1 Proofs of (39) and (40)
We provide the complete derivation for the upper bounds on the expected regret terms $\mathbb{E}\{R_{1}\}$, $\mathbb{E}\{R_{2}\}$, and $\mathbb{E}\{R_{3}\}$ for the Blocked Hedge+OGD algorithm with block size $H$. Let $M=\lceil T/H\rceil$ be the total number of blocks. We denote the blocks by $\mathcal{B}_{1},\dots,\mathcal{B}_{M}$.
O.1.1 Upper Bound for $\mathbb{E}\{R_{1}\}$
The term $R_{1}$ captures the regret of the weight updates $\theta_{m}$ against the best fixed distribution $\theta^{\star}$ over the sequence of blocks:
$R_{1}=\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}\Lambda^{(m)}(x_{m}),$
where $\Lambda^{(m)}(x)=\sum_{t\in\mathcal{B}_{m}}\lambda_{t}(x)$ is the cumulative loss vector for block $m$.
Proof.
  1. 1.
    The loss vector for each step $m$ is the block sum $\Lambda^{(m)}(x_{m})$. Given that $\|\lambda_{t}\|_{\infty}\leq B$ (Assumption 2), the magnitude of the loss vector is bounded by:
    $\|\Lambda^{(m)}(x_{m})\|_{\infty}\leq\sum_{t\in\mathcal{B}_{m}}\|\lambda_{t}\|_{\infty}\leq H\cdot B\text{.}$
  2. 2.
    Although Algorithm 5 updates weights using the block-average loss $\bar{\lambda}_{m}=\Lambda^{(m)}/H_{m}$, this is mathematically equivalent to the gains version of Hedge running on the meta-sequence of blocks with an effective step size $\tilde{\eta}_{\theta,m}=\eta_{\theta,m}/H$.
  3. 3.
    For a horizon of $M$ steps with losses bounded by $L_{\max}=HB$ and $K=2$ experts, the Hedge algorithm yields:
    $R_{1}\leq HB\sqrt{2M\ln 2}\text{.}$
  4. 4.
    Using $M\approx T/H$
    $R_{1}\leq HB\sqrt{\frac{2T\ln 2}{H}}=B\sqrt{2TH\ln 2}\text{.}$
Thus, $\mathbb{E}\{R_{1}\}=O(\sqrt{TH})$. ∎
O.1.2 Upper Bound for $\mathbb{E}\{R_{2}\}$
The term $R_{2}$ captures the regret of the action updates $x_{m}$ against the best fixed action $x^{\star}$:
$R_{2}=\sum_{m=1}^{M}\left(\sum_{t\in\mathcal{B}_{m}}\theta_{m}^{\top}\lambda_{t}(x_{m})-\sum_{t\in\mathcal{B}_{m}}\theta_{m}^{\top}\lambda_{t}(x^{\star})\right)\text{.}$
Proof.
  1. 1.
    The effective gradient for block $m$ is the sum of gradients: $G^{(m)}=\sum_{t\in\mathcal{B}_{m}}\nabla(\theta_{m}^{\top}\lambda_{t}(x_{m}))$. Given $\|\nabla\lambda_{t}\|\leq G$ (Assumption 3), the gradient norm is bounded by:
    $\|G^{(m)}\|\leq\sum_{t\in\mathcal{B}_{m}}\|\nabla\lambda_{t}\|\leq H\cdot G\text{.}$
  2. 2.
    Algorithm 5 performs OGD using the block-average gradient $\bar{g}_{m}=G^{(m)}/H_{m}$. In the regret analysis, this is equivalent to OGD using the cumulative gradient $G^{(m)}$ with an effective step size $\tilde{\eta}_{x,m}=\eta_{x,m}/H$.
  3. 3.
    For $M$ steps with gradient norms bounded by $\mathcal{G}_{\max}=HG$ and domain diameter $D$, OGD yields:
    $R_{2}\leq D(HG)\sqrt{M}=DHG\sqrt{\frac{T}{H}}=DG\sqrt{TH}\text{.}$
Thus, $\mathbb{E}\{R_{2}\}=O(\sqrt{TH})$. ∎
O.1.3 Upper Bound for $\mathbb{E}\{R_{3}\}$
$\begin{aligned} R_{3}&=\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x)-{\bar{C}}_{\mathsf{OPT}},\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x^{\star})-\sum_{m=1}^{M}H(\theta^{\star})^{\top}\mu(x^{\star}),\\ &\stackrel{{\scriptstyle(b)}}{{=}}\sum_{m=1}^{M}\theta_{m}^{\top}(\Lambda^{(m)}(x^{\star})-H\mu(x^{\star}))+\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star})),\\ &\stackrel{{\scriptstyle(c)}}{{=}}\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(\Lambda^{(m)}(x^{\star})-H\mu(x^{\star}))+\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star}))\\ &\quad+\sum_{m=1}^{M}(\theta^{\star})^{\top}(\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})),\\ &\stackrel{{\scriptstyle(d)}}{{\leq}}\sum_{m=1}^{M}\|\theta_{m}-\theta^{\star}\|_{1}\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty}+\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star}))\\ &\quad+\sum_{m=1}^{M}\|\theta^{\star}\|_{1}\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty},\\ &\stackrel{{\scriptstyle(e)}}{{\leq}}3\sum_{m=1}^{M}\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty},\end{aligned},$
where $(a)$ follows from the suboptimality of the stationary optimal action $x^{\star}\in\mathcal{X}$, $(b)$ is obtained by adding and subtracting the mixed term $\sum_{m=1}^{M}\theta_{m}^{\top}(H\mu(x^{\star}))$, $(c)$ uses the algebraic identity $\theta_{m}^{\top}V=(\theta_{m}-\theta^{\star})^{\top}V+(\theta^{\star})^{\top}V$, where $V=\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})$, to center the algorithm’s weights around the stationary optimal weights $\theta^{\star}$, $(d)$ follows from Holder’s inequality, where $u^{\top}v\leq\|u\|_{1}\|v\|_{\infty}$, and finally $(e)$ utilizes three key properties:
  1. 1.
    since $\theta_{m},\theta^{\star}\in\Delta_{K}$, we have $\|\theta_{m}-\theta^{\star}\|_{1}\leq\|\theta_{m}\|_{1}+\|\theta^{\star}\|_{1}=2$,
  2. 2.
    stationary optimality of $\theta^{\star}$ ensures $(\theta_{m}-\theta^{\star})^{\top}\mu(x^{\star})\leq 0$ for all $\theta_{m}\in\Delta_{K}$, so $\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star}))$ is bounded by zero, and
  3. 3.
    $\|\theta^{\star}\|_{1}=1$ because it is a probability distribution on the simplex.
To complete the proof, we now bound the expected deviation term $\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty}$ using the mixing properties of the chain. By the law of iterated expectations, we have $\mathbb{E}[\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty}]=\mathbb{E}[\mathbb{E}[\|\sum_{t\in\mathcal{B}_{m}}(\lambda_{t}(x^{\star})-\mu(x^{\star}))\|_{\infty}\mid\mathcal{F}_{\text{start}(m)}]]$, where $\mathcal{F}_{\text{start}(m)}$ is the filtration at the start of block $m$.
  1. 1.
    For a uniformly ergodic Markov chain, the expected deviation of the state from stationarity decays exponentially with time. Let $S_{\text{start}(m)}$ be the state at the start of block $m$. For the $k$-th step within the block ($1\leq k\leq H$):
    $\left\|\mathbb{E}[\lambda_{(m-1)H+k}(x^{\star})\mid S_{\text{start}(m)}]-\mu(x^{\star})\right\|_{\infty}\leq Ce^{-k/\tau_{\mathrm{mix}}}$
  2. 2.
    We sum this deviation over the $H$ steps in the block:
    $\left\|\mathbb{E}\left[\sum_{t\in\mathcal{B}_{m}}(\lambda_{t}(x^{\star})-\mu(x^{\star}))\;\Bigg|\;\mathcal{F}_{\text{start}(m)}\right]\right\|_{\infty}\leq\sum_{k=1}^{H}Ce^{-k/\tau_{\mathrm{mix}}}.$
    This geometric series is bounded by a constant proportional to the mixing time
    $\sum_{k=1}^{\infty}Ce^{-k/\tau_{\mathrm{mix}}}=O(\tau_{\mathrm{mix}})$
    Thus, the total bias accumulated per block is $O(\tau_{\mathrm{mix}})$.
  3. 3.
    Summing this constant error over the $M$ blocks
    $\mathbb{E}\{R_{3}\}\leq 3\sum_{m=1}^{M}O(\tau_{\mathrm{mix}})=M\cdot O(\tau_{\mathrm{mix}})=\frac{T}{H}O(\tau_{\mathrm{mix}})$
Thus, we get
$\mathbb{E}\{R_{3}\}=O\left(\frac{T\tau_{\mathrm{mix}}}{H}\right).$